# StableMotion: Training Motion Cleanup Models with Unpaired Corrupted Data

### v1

@@ Abstract.
Motion capture (mocap) data often exhibits visually jarring artifacts due to inaccurate sensors and post-processing. Cleaning this corrupted data can require substantial manual effort from human experts, which can be a costly and time-consuming process. Previous data-driven motion cleanup methods offer the promise of automating this cleanup process, but often require in-domain paired corrupted-to-clean training data. Constructing such paired datasets requires access to high-quality, relatively artifact-free motion clips, which often necessitates laborious manual cleanup. In this work, we present StableMotion, a simple yet effective method for training motion cleanup models directly from unpaired corrupted datasets that need cleanup. The core component of our method is the introduction of motion quality indicators, which can be easily annotated— through manual labeling or heuristic algorithms—and enable training of quality-aware motion generation models on raw motion data with mixed quality. At test time, the model can be prompted to generate high-quality motions using the quality indicators. Our method can be implemented through a simple diffusion-based framework, leading to a unified motion generate-discriminate model, which can be used to both identify and fix corrupted frames. We demonstrate that our proposed method is effective for training motion cleanup models on raw mocap data in production scenarios by applying StableMotion to SoccerMocap, a 245-hour soccer mocap dataset containing real-world motion artifacts. The trained model effectively corrects a wide range of motion artifacts, reducing motion pops and frozen frames by 68% and 81%, respectively. On our benchmark dataset, we further show that cleanup models trained with our method on unpaired corrupted data outperform state-of-the-art methods trained on clean or paired data, while also achieving comparable performance in preserving the content of the original motion clips. Code and benchmark models will be released upon publication of this paper. See https://youtu.be/3Y7MMAH02B4 for more results.
@@
ccs: Computing methodologies Motion processing
Figure 1. StableMotion enables effective training of motion cleanup models using raw mocap data containing real-world artifacts, without requiring paired corrupted-to-clean motion data. Once trained, the model can automatically identify and fix corrupted motions across a wide range of real-world artifact types. This example showcases automatic correction of body self-penetration and unnatural wrist movements.
@@ 1. Introduction
High-quality motion data serves as a cornerstone for numerous applications in computer animation. With the increasing prevalence of deep learning techniques for motion synthesis, the demand for large-scale, high-quality motion datasets has grown significantly (Guo et al., 2022; Peng et al., 2022; Shi et al., 2024; Tevet et al., 2023; Alexanderson et al., 2023). However, such datasets are often scarce and challenging to procure. Even motion data captured by professional mocap studios often contain artifacts caused by occlusion, sensor inaccuracies, or errors in post-processing (Holden, 2018). These issues are further exacerbated when collecting motion data from in-the-wild environments or outdoor scenarios, which are common sources for large-scale datasets (Lin et al., 2024; Jiang et al., 2024). Cleaning noisy motion data requires significant manual effort from artists, resulting in a costly and labor-intensive process, especially in large-scale production environments that rely on vast amounts of mocap data. Data-driven motion cleanup models present a compelling solution by automating the cleanup of noisy data using learned priors from large motion datasets. However, most existing methods for developing motion cleanup models depend on paired datasets of clean and corrupted motions, which can be used to train supervised learning models to correct motion artifacts (Zhang et al., 2024; Li et al., 2025; Holden, 2018; Chen et al., 2021b). The creation of such paired training data is inherently challenging, requiring either costly manual cleanup of corrupted data or augmentation with synthetic artifacts on already-cleaned data, which must accurately replicate real-world artifacts. These challenge becomes especially significant for large-scale, domain-specific applications, such as sports.
@@
A similar challenge exists in reinforcement learning (RL), where it is much more costly to collect expert demonstrations compared to suboptimal demonstrations for a specific task. Surprisingly, Kumar et al. (2022) observed that offline RL with suboptimal data could surpass behavior cloning (BC) with expert-only data. For instance, Decision Transformer (Chen et al., 2021a), which models the state-return trajectory, are able to generate novel optimal solutions for finding shortest path in a graph and outperforms BC on Atari games by leveraging diverse suboptimal data. This performance improvement may stem from the richer diversity inherent in suboptimal data, which enables the policy to learn not only the desired behaviors for a given task, but also which behaviors to avoid (Kumar et al., 2022). Once the model has been trained, the desired performance of the policy can subsequently be specified by conditioning the policy on the appropriate target return. Inspired by state-return trajectory modeling in offline RL, we incorporate a frame-level quality indicator variable (QualVar) similar to state-level reward in RL. Our StableMotion framework utilizes a generate-discriminate approach, where a model is jointly trained to evaluate motion quality and generate motion of varying quality, according to QualVar. Similar to the practice of prompting text-to-image model with “photo-realistic quality”, QualVar offers a knob to specify the generation quality. We further explicitly integrate quality-conditioned motion generation and motion-conditioned QualVar prediction into a diffusion-based model. Once trained, the model can: 1) evaluate the quality of a given motion, and 2) generate high-quality motions by conditioning on the desired quality variables. This allows the model to cleanup raw mocap data by first identifying corrupted segments and then inpainting them with high-quality motions.
@@
The central contribution of this work is StableMotion, a framework for training motion cleanup models from unpaired, corrupted motion data. Instead of learning to directly restore corrupted motions to clean motions, our method leverages all available raw motion data, with mixed quality, using a simple conditional inpainting framework. The trained generate-discriminate model is capable of both evaluating motion quality and generating higher-quality motions than those in the original training dataset. Our method can fix corrupted segments by generating higher-quality alternatives, while leaving uncorrupted frames intact. We also propose a number of test-time techniques that harness the sample diversity and the dual functionality of generate-discriminate model to improve consistency and preservation of the content in the original motion clips. We apply StableMotion to train a cleanup model on a 245-hour, large-scale soccer mocap dataset, SoccerMocap, containing artifacts from real-world capture systems. The model successfully corrects a wide range of motion artifacts, reducing motion pops and frozen frames by 68% and 81%, respectively, showcasing StableMotion’s efficacy in real-world production scenarios.
@@ 2. Related Works
Another common approach for motion cleanup involves kinematic-based data-driven methods. Many kinematic-based regression methods have been proposed to restore clean motions from their corrupted counterparts (Mall et al., 2017; Holden, 2018; Pan et al., 2024, 2023; Chen et al., 2021b; Zhang et al., 2024; Li et al., 2025), achieving state-of-the-art performance on motion cleanup tasks. However, these methods require paired corrupted-to-clean data for training, which is often impractical in real production scenarios. Such paired data are often created by synthetically corrupting already-cleaned general-purpose datasets, such as AMASS and LAFAN1 (Mahmood et al., 2019; Harvey et al., 2020), which requires emulating the characteristics of real-world artifacts. To bypass this challenge, previous methods have also trained models solely on clean datasets and taken advantage of the regularization characteristics of data-driven frameworks to mitigate abnormal behaviors. For instance, autoencoding frameworks can be used to learn compact latent representations of the underlying motion manifolds (Holden et al., 2016, 2015; Aberman et al., 2020). Once the manifold has been constructed, motion artifacts can be fixed by encoding and decoding through the latent space, leveraging the structure of the latent manifold to regularize and remove artifacts. Similarly, motion synthesis models have also demonstrated promising results for motion cleanup by inpainting corrupted frames with high-quality counterparts using generative models trained on large general-purpose datasets (Cohan et al., 2024; Shafir et al., 2024; Tevet et al., 2023; Guo et al., 2024). Recent works that combine the advantages of inpainting and the denoising nature of diffusion models have shown more promising results in fixing artifacts while preserving the original content (Hwang et al., 2025), demonstrating the state-of-the-art performance in post-processing noisy video-based mocap. However, one of the common limitations of these methods is the need for large-scale, high-quality datasets to learn an approximation of the manifolds of natural motions. A model trained on a general-purpose motion dataset may not generalize effectively when applied to domain-specific data, such as sport-related mocap, which are common in production scenarios. On the other hand, acquiring a high-quality in-domain dataset for training domain-specific cleanup models can also be prohibitively expensive, often requiring manual cleanup, and thereby limiting these methods’ real-world applicability.
@@ 3. Background: Motion Diffusion Model
To approximate the inverse diffusion process for generation, we follow prior motion diffusion models and train a denoising model $g$ to predict the fully denoised sample ${\mathbf{x}}_{0}$ (Tevet et al., 2023). The denoising model $g(\hat{{\mathbf{x}}}_{0}|{\mathbf{x}}_{t})$ is trained using the simple DDPM objective (Ho et al., 2020):
(2) $\mathcal{L}_{\text{simple}}=\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{M}},t\sim\mathcal{U}[1,T],{\mathbf{x}}_{t}\sim q({\mathbf{x}}_{t}|{\mathbf{x}}_{0})}\left[\|{\mathbf{x}}_{0}-g(\hat{{\mathbf{x}}}_{0}|{\mathbf{x}}_{t})\|^{2}\right].$
@@ 4. Overview
In this work, our primary goal is to develop a framework for training motion cleanup models directly on raw, domain-specific motion datasets. These models can then be applied to clean artifacts in the desired dataset, resulting in high-quality motion data tailored to the specific domain. Models trained on general-purpose datasets, such as AMASS (Mahmood et al., 2019), often fail to generalize effectively to domain-specific data due to differences in depicted behaviors and retargeting errors arising from skeleton mismatches. For instance, the dynamic movements of athletes in sports mocap datasets differ significantly from the everyday motions typically found in general-purpose datasets. In such applications, training cleanup models directly on domain-specific data is vital for effective motion cleanup. However, real-world motion capture data often contain artifacts (Holden, 2018), which may be randomly distributed throughout motion clips. Constructing paired corrupted-to-clean data for supervised training is both expensive and labor-intensive. Moreover, naively discarding corrupted segments fragments the dataset into short, disconnected clips (e.g., leading to an average clip length of approximately 1.5s in SoccerMocap), undermining the temporal coherence required for learning long-horizon motion dynamics. These challenges highlight the need for a data-driven framework that can be trained directly on raw mocap data, without breaking the continuity of the underlying sequences.
An overview of StableMotion is illustrated in Figure 2. Our generate-discriminate framework unifies the tasks of motion quality evaluation and quality-aware motion generation through a simple diffusion objective (Section 5). During training, the model receives as input a raw motion sequence ${\mathbf{m}}^{1:N}$, along with quality indicators ${\mathbf{h}}^{1:N}$ for each frame, where corrupted frames are marked as ${\mathbf{h}}^{i}=\mathbf{1}$ and clean frames as ${\mathbf{h}}^{i}=\mathbf{0}$. The model is trained to jointly generate the motion ${\mathbf{m}}$ of mixed quality and the corresponding QualVar ${\mathbf{h}}$ through a diffusion-denoising process. The trained model can be used to automatically clean motion data by first predicting QualVar to identify corrupted frames, and then inpainting corrupted frames with higher-quality motions (Section 6). We propose a number of additional test-time techniques to further improve consistency and preserve the content of the original motion clips.
@@ 5. Generate-Discriminate Model
The core component of our framework is a generate-discriminate model, which generates content alongside its corresponding discriminative labels. In this work, the discriminative labels are per-frame quality indicator variables (QualVar), which enable training quality-aware models from unpaired, corrupted data of mixed quality. The indicator variables also provide a mechanism to direct the model to produce clean motion by conditioning on desired QualVar. In this section, we first present the process of training quality-aware diffusion models using unpaired, corrupted motion data. Then, we detail the test-time process of applying the model to identify and clean motion artifacts.
@@ 5.1. Training: Quality-Aware Inpainting
The standard motion diffusion model provides a strong model for generating motions that follow a given dataset distribution (Tevet et al., 2023), but it cannot selectively generate motions of a desired quality when the dataset contains artifacts. To directly train a motion cleanup model on uncleaned raw mocap data, we introduce a QualVar, ${\mathbf{h}}$, to represent the underlying distribution of different data quality levels. We train our model directly on the combined features of the motion and its corresponding QualVar, ${\mathbf{x}}_{t}=({\mathbf{m}},{\mathbf{h}})_{t}$, where $t$ denotes the diffusion step. By incorporating QualVar into the feature representation, the motion diffusion model can represent motions of varying quality with explicit control over the quality of the generated motions. This simple approach enables effective training of motion diffusion models on raw corrupted motion data, which can then generate high-quality clean motion.
To further tailor the model for motion cleanup, we design two inpainting training tasks to facilitate the key functions of our motion cleanup model: 1) motion quality evaluation and 2) quality-aware motion generation. At runtime, the model is provided with either motion frames or QualVar, and the model then infers the missing features. While prior methods have shown that an unconditional motion diffusion model $g(\hat{{\mathbf{x}}}_{0}|{\mathbf{x}}_{t})$ can perform similar inpainting tasks using inference-time imputation techniques (Lugmayr et al., 2022; Tevet et al., 2023), recent work has demonstrated that explicitly training a conditional inpainting model tends to outperform methods relying solely on test-time imputation strategies with an unconditional model (Cohan et al., 2024; Karunratanakul et al., 2023). Based on these observations, we adopt a conditional model for motion cleanup. Given an inpainting mask ${\mathbf{b}}$, where observed conditional features are marked as 1 and inpainting features as 0, a quality-aware motion diffusion model can be trained using a masked inpainting objective. The mask is randomly generated in two modes: 1) motion evaluation and 2) motion generation. For motion evaluation, the goal is to infer the quality indicator variables ${\mathbf{h}}^{1:N}$ for all frames, given the motion sequence ${\mathbf{m}}^{1:N}$. For motion generation, we denote the missing frame indices as $i\in{\mathbb{A}}\subseteq\{1,\dots,N\}$ and the observed frame indices as $i\in{\mathbb{B}}={\mathbb{A}}^{\complement}$. The goal is to generate the missing motion frames ${\mathbf{m}}^{i\in{\mathbb{A}}}$ given the observed frames ${\mathbf{m}}^{i\in{\mathbb{B}}}$ and the per-frame specified quality indicators ${\mathbf{h}}^{1:N}$. Both modes can be easily implemented through different inpainting masks, providing a unified quality-aware inpainting objective.

### v2

@@ Abstract.
Motion capture (mocap) data often exhibits visually jarring artifacts due to inaccurate sensors and post-processing. Cleaning this corrupted data can require substantial manual effort from human experts, which can be a costly and time-consuming process. Previous data-driven motion cleanup methods offer the promise of automating this cleanup process, but often require in-domain paired corrupted-to-clean training data. Constructing such paired datasets requires access to high-quality, relatively artifact-free motion clips, which often necessitates laborious manual cleanup. In this work, we present StableMotion, a simple yet effective method for training motion cleanup models directly from unpaired corrupted datasets that need cleanup. The core component of our method is the introduction of motion quality indicators, which can be easily annotated— through manual labeling or heuristic algorithms—and enable training of quality-aware motion generation models on raw motion data with mixed quality. At test time, the model can be prompted to generate high-quality motions using the quality indicators. Our method can be implemented through a simple diffusion-based framework, leading to a unified motion generate-discriminate model, which can be used to both identify and fix corrupted frames. We demonstrate that our proposed method is effective for training motion cleanup models on raw mocap data in production scenarios by applying StableMotion to SoccerMocap, a 245-hour soccer mocap dataset containing real-world motion artifacts. The trained model effectively corrects a wide range of motion artifacts, reducing motion pops and frozen frames by 68% and 81%, respectively. On our benchmark dataset, we further show that cleanup models trained with our method on unpaired corrupted data outperform state-of-the-art methods trained on clean or paired data, while also achieving comparable performance in preserving the content of the original motion clips.
@@
copyright: acmcopyrightcopyright: acmlicensedcopyright: rightsretainedjournalyear: 2025conference: SIGGRAPH Asia 2025 Conference Papers; December 15–18, 2025; Hong Kong, Hong Kongbooktitle: SIGGRAPH Asia 2025 Conference Papers (SA Conference Papers ’25), December 15–18, 2025, Hong Kong, Hong Kongdoi: 10.1145/3757377.3763817isbn: 979-8-4007-2137-3/2025/12ccs: Computing methodologies Motion processing
Figure 1. StableMotion enables effective training of motion cleanup models using raw mocap data containing real-world artifacts, without requiring paired corrupted-to-clean motion data. Once trained, the model can automatically identify and fix corrupted motions across a wide range of real-world artifacts. This example showcases automatic correction of body self-penetration and unnatural wrist movements.
@@ 1. Introduction
High-quality motion data serves as a cornerstone for numerous applications in computer animation. With the increasing prevalence of deep learning techniques for motion synthesis, the demand for large-scale, high-quality motion datasets has grown significantly (Guo et al., 2022; Peng et al., 2022; Shi et al., 2024; Tevet et al., 2023; Alexanderson et al., 2023). However, such datasets are often scarce and challenging to procure. Even motion data captured by professional mocap studios often contain artifacts caused by occlusion, sensor inaccuracies, or errors in post-processing (Holden, 2018). These issues are further exacerbated when collecting motion data from in-the-wild environments or outdoor scenarios, which are common sources for large-scale datasets (Lin et al., 2024; Jiang et al., 2024). Cleaning noisy motion data requires significant manual effort from artists, resulting in a costly and labor-intensive process, especially in large-scale production environments that rely on vast amounts of mocap data. Data-driven motion cleanup models present a compelling solution by automating the cleanup of noisy data using learned priors from large motion datasets. However, most existing methods for developing motion cleanup models depend on paired datasets of clean and corrupted motions, which are then used to train supervised learning models to correct motion artifacts (Zhang et al., 2024; Li et al., 2025; Holden, 2018; Chen et al., 2021b). The creation of such paired training data is inherently challenging, requiring either costly manual cleanup of corrupted data or augmentation with synthetic artifacts on already-cleaned data, which must accurately replicate real-world artifacts. These challenges become especially daunting for large-scale, domain-specific applications, such as sports, where the cost of manual cleanup of these large datasets can quickly become exorbitant.
@@
A similar challenge exists in reinforcement learning (RL), where it is much more costly to collect expert demonstrations compared to suboptimal demonstrations for a specific task. Surprisingly, Kumar et al. (2022) observed that offline RL with suboptimal data could surpass behavior cloning (BC) with expert-only data. For instance, Decision Transformer (Chen et al., 2021a), which models the state-return trajectories, are able to generate novel optimal solutions for finding shortest paths in a graph and outperforms BC on Atari games by leveraging diverse suboptimal data. This performance improvement may stem from the richer diversity inherent in suboptimal data, which enables the policy to learn not only the desired behaviors for a given task, but also which behaviors to avoid (Kumar et al., 2022). Once the model has been trained, the desired performance of the policy can subsequently be generated by conditioning the policy on the appropriate target return. Inspired by state-return trajectory modeling in offline RL, we incorporate a frame-level quality indicator variable (QualVar) similar to state-level reward in RL. Our StableMotion framework utilizes a generate-discriminate approach, where a model is jointly trained to evaluate motion quality and generate motion of varying quality, according to QualVar. Similar to the practice of prompting text-to-image model with “photo-realistic quality”, QualVar offers a knob to specify the generation quality. We further explicitly integrate quality-conditioned motion generation and motion-conditioned QualVar prediction into a unified diffusion-based model. Once trained, the model can: 1) evaluate the quality of a given motion, and 2) generate high-quality motions by conditioning on the desired quality variables. This allows the model to cleanup raw mocap data by first identifying corrupted segments and then inpainting them with high-quality motions.
@@
The central contribution of this work is StableMotion, a framework for training motion cleanup models from unpaired, corrupted motion data. Instead of learning to directly restore corrupted motions to clean motions, our method leverages all available raw motion data, with mixed quality, using a simple conditional inpainting framework. The trained generate-discriminate model is capable of both evaluating motion quality and generating higher quality motions than those in the original training dataset. Our method can fix corrupted segments by generating higher-quality alternatives, while leaving uncorrupted frames intact. We also propose a number of test-time techniques that harness the sample diversity and the dual functionality of the generate-discriminate model to improve consistency and preservation of the content in the original motion clips. We apply StableMotion to train a cleanup model on a 245-hour, large-scale soccer mocap dataset, SoccerMocap, containing artifacts from real-world capture systems. The model successfully corrects a wide range of motion artifacts, reducing motion pops and frozen frames by 68% and 81%, respectively, showcasing StableMotion’s efficacy in real-world production scenarios.
@@ 2. Related Works
Another common approach for motion cleanup involves kinematic-based data-driven methods. Many kinematic-based regression methods have been proposed to restore clean motions from their corrupted counterparts (Mall et al., 2017; Holden, 2018; Pan et al., 2024, 2023; Chen et al., 2021b; Zhang et al., 2024; Li et al., 2025), achieving state-of-the-art performance on motion cleanup tasks. However, these methods require paired corrupted-to-clean data for training, which is often impractical in real production scenarios. Such paired data are often created by synthetically corrupting already-cleaned general-purpose datasets, such as AMASS and LAFAN1 (Mahmood et al., 2019; Harvey et al., 2020), which requires emulating the characteristics of real-world artifacts. To bypass this challenge, previous methods have also trained models solely on clean datasets and taken advantage of the regularization characteristics of data-driven frameworks to mitigate abnormal behaviors. For instance, autoencoding frameworks can be used to learn compact latent representations of the underlying motion manifolds (Holden et al., 2016, 2015; Aberman et al., 2020). Once the manifold has been constructed, motion artifacts can be fixed by encoding and decoding through the latent space, leveraging the structure of the latent manifold to regularize and remove artifacts. Similarly, motion synthesis models have also demonstrated promising results for motion cleanup by inpainting corrupted frames with high-quality counterparts using generative models trained on large general-purpose datasets (Cohan et al., 2024; Shafir et al., 2024; Tevet et al., 2023; Guo et al., 2024). Recent works that combine the advantages of inpainting and the denoising nature of diffusion models have shown more promising results in fixing artifacts while preserving the original content (Hwang et al., 2025), demonstrating the state-of-the-art performance in post-processing noisy video-based mocap. However, a common limitation of these methods is the need for large-scale, high-quality datasets in order to train models to approximate the manifolds of natural motions (Harvey and Pal, 2018; Harvey et al., 2020; Hernandez et al., 2019; Kaufmann et al., 2020). A model trained on a general-purpose motion dataset may not generalize effectively when applied to domain-specific data, such as sport-related mocap, which are common in production scenarios. On the other hand, acquiring a high-quality in-domain dataset for training domain-specific cleanup models can also be prohibitively expensive, often requiring manual cleanup, and thereby limiting these methods’ real-world applicability.
@@ 3. Background: Motion Diffusion Model
To approximate the inverse diffusion process for generation, we follow prior motion diffusion models and train a denoising model $g$ to predict the fully denoised sample ${\mathbf{x}}_{0}$ (Tevet et al., 2023). The denoising model $\hat{{\mathbf{x}}}_{0}=g({\mathbf{x}}_{t})$ is trained using the simple DDPM objective (Ho et al., 2020):
(2) $\mathcal{L}_{\text{simple}}=\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{M}},t\sim\mathcal{U}[1,T],{\mathbf{x}}_{t}\sim q({\mathbf{x}}_{t}|{\mathbf{x}}_{0})}\left[\|{\mathbf{x}}_{0}-g({\mathbf{x}}_{t})\|^{2}\right].$
@@ 4. Overview
In this work, our primary goal is to develop a framework for training motion cleanup models directly on raw, domain-specific motion datasets. These models can then be applied to clean artifacts in the desired dataset, resulting in high-quality motion data tailored to the specific domain. Models trained on general-purpose datasets, such as AMASS (Mahmood et al., 2019), often fail to generalize effectively to domain-specific data due to differences in depicted behaviors and retargeting errors arising from morphological differences. For instance, the dynamic movements of athletes in sports mocap datasets differ significantly from the everyday motions typically found in general-purpose datasets. In such applications, training cleanup models directly on domain-specific data is vital for effective motion cleanup. However, real-world motion capture data often contain artifacts (Holden, 2018), which may be randomly distributed throughout motion clips. Constructing paired corrupted-to-clean data for supervised training is both expensive and labor-intensive. Moreover, naively discarding corrupted segments will fragment the dataset into short, disconnected clips (e.g., leading to an average clip length of approximately 1.5s in SoccerMocap), undermining the temporal coherence required for learning long-horizon motion dynamics (see Table 1 w/ Filtered). These challenges highlight the need for a data-driven framework that can be trained directly on raw mocap data, without breaking the continuity of the underlying sequences.
An overview of StableMotion is illustrated in Figure 2. Our generate-discriminate framework unifies the tasks of motion quality evaluation and quality-aware motion generation through a simple inpainting objective (Section 5). During training, the model receives as input a raw motion sequence ${\mathbf{m}}^{1:N}$, along with quality indicators ${\mathbf{h}}^{1:N}$ for each frame, where corrupted frames are marked as ${\mathbf{h}}^{i}=\mathbf{1}$ and clean frames as ${\mathbf{h}}^{i}=\mathbf{0}$. The model is trained to jointly generate the motion $\mathbf{m}$} of mixed quality and the corresponding QualVar $\mathbf{h}$} through a diffusion-denoising process. The trained model can be used to automatically clean motion data by first predicting QualVar to identify corrupted frames, and then inpainting corrupted frames with higher-quality motions (Section 6). We propose a number of additional test-time techniques to further improve consistency and preserve the content of the original motion clips.
@@ 5. Generate-Discriminate Model
The core component of our framework is a generate-discriminate model, which generates motions alongside their corresponding discriminative labels. In this work, the discriminative labels are per-frame quality indicator variables (QualVar), which enable training quality-aware models from unpaired, corrupted data of mixed quality. The indicator variables also provide a mechanism to direct the model to produce clean motions by conditioning on desired the QualVar. In this section, we first present the training process of quality-aware diffusion models using unpaired, corrupted motion data. Then, we detail the test-time process of applying the model to identify and clean motion artifacts.
@@ 5.1. Training: Quality-Aware Inpainting
The standard motion diffusion model provides a strong model for generating motions that follow a given dataset distribution (Tevet et al., 2023), but it cannot selectively generate motions of a desired quality when the dataset contains artifacts. To directly train a motion cleanup model on uncleaned raw mocap data, we introduce quality indicator variables $\mathbf{h}$}, to represent the underlying distribution of different data quality levels. We train our model directly on the combined features of the motion and its corresponding QualVar, ${\mathbf{x}}_{t}=({\mathbf{m}},{\mathbf{h}})_{t}$, where $t$ denotes the diffusion step. By incorporating QualVar into the feature representation, the motion diffusion model can represent motions of varying quality with explicit control over the quality of the generated motions. This simple approach enables effective training of motion diffusion models on raw corrupted motion data, which can then generate high-quality clean motion.
To further tailor the model for motion cleanup, we design two inpainting training tasks to facilitate the key functions of our motion cleanup model: 1) motion quality evaluation and 2) quality-aware motion generation. At runtime, the model is provided with either motion frames or QualVar, and the model then infers the missing features. While prior methods have shown that an unconditional motion diffusion model $g({\mathbf{x}}_{t})$ can perform similar inpainting tasks using inference-time imputation techniques (Lugmayr et al., 2022; Tevet et al., 2023), recent work has demonstrated that explicitly training a conditional inpainting model tends to outperform methods relying solely on test-time imputation strategies with an unconditional model (Cohan et al., 2024; Karunratanakul et al., 2023). Based on these observations, we adopt a conditional model for motion cleanup. Given an inpainting mask $\mathbf{b}$}, where observed conditional features are marked as 1 and inpainting features as 0, a quality-aware motion diffusion model can be trained using a masked inpainting objective. The mask is randomly generated in two modes: 1) motion evaluation and 2) motion generation. For motion evaluation, the goal is to infer the quality indicator variables ${\mathbf{h}}^{1:N}$ for all frames, given the motion sequence ${\mathbf{m}}^{1:N}$. For motion generation, we denote the missing frame indices as $i\in{\mathbb{A}}\subseteq\{1,\dots,N\}$ and the observed frame indices as $i\in{\mathbb{B}}={\mathbb{A}}^{\complement}$. The goal is to generate the missing motion frames ${\mathbf{m}}^{i\in{\mathbb{A}}}$ given the observed frames ${\mathbf{m}}^{i\in{\mathbb{B}}}$ and the per-frame specified quality indicators ${\mathbf{h}}^{1:N}$. Both modes can be easily implemented through different inpainting masks, providing a unified quality-aware inpainting objective.